OpenStudy (anonymous):

Now I Need Help On This One ... Can There Be A Square Root For -147? I Got 12.12....

OpenStudy (lalaly):

no its complex... so its 12.12i

OpenStudy (anonymous):

\[\sqrt{-1}=i\]

OpenStudy (lalaly):

http://www.wolframalpha.com/input/?i=sqrt+-147

OpenStudy (anonymous):

You take out the -1 first then work on the rest.

OpenStudy (anonymous):

\[\sqrt{-1*147}=i\sqrt{147}\]

OpenStudy (anonymous):

so, like there is two blank spaces and theres a square root 3